A beam of protons are all traveling in the same dilooking for help on all three portions including the bonus questios.
Answer #1

a) Since the proton has a positive charge and the bending of motion is clockwise hence the direction of the magnetic field will, be out of page direction so that the velocity vector cross product with magnetic flux density vector could be towards the right at each point.

b) Here the radius of the circular motion is 15 cm and the magnitude of magnetic field is 2.5 mT then the formula:

q B

where rho is the radius , m is mass, v is velocity, q is charge, B is magnetic field.

0.15 = 1.67*10^-27*v/(1.6*10^-19*2.5*10^-3)

v = (1.6*10^-19*2.5*10^-3)*0.15/(1.67*10^-27) = 35928.1437 m/s